# Roadmaps with Gaps over Controllers: Achieving Efficiency in Planning under Dynamics

### v3

@@ V Conclusion
This paper proposes a strategy that can benefit from learned controllers to improve the efficiency of kinodynamic planning for robots with significant dynamics. It utilizes a controller trained offline for a robot in an empty environment. The target environment is represented via a “Roadmap with Gaps” over local regions and applications of the controller between them. Given a wavefront over the roadmap for a specific goal, a tree sampling-based motion planner generates informed subgoals and uses the controller to reach them. When the controller cannot reach a subgoal, the planner resorts to random exploration. Experiments on various benchmarks demonstrate significant improvement in planning efficiency.
For higher-dimensional systems, the memory requirements of the roadmap can be improved by considering sparse representations similar to geometric planning [36]. Furthermore, learned reachability estimators can assist in efficient roadmap construction and online queries. This work assumes that an accurate model of the environment and the vehicular system is available, which complicates the direct deployment of the proposed approach on real vehicular systems. This also motivates integrating the proposed motion planner with system identification tools as well as state estimation and feedback control to track the planned trajectory.

### v5

@@ V Conclusion
This paper proposes a strategy that can benefit from learned controllers to improve the efficiency of kinodynamic planning for robots with significant dynamics. It utilizes a controller trained offline in an empty environment. The target environment is represented via a “Roadmap with Gaps” over local regions and applications of the controller between them. Given a wavefront over the roadmap for a specific goal, a tree sampling-based motion planner generates informed subgoals and uses the controller to reach them. When the controller cannot reach a subgoal, the planner resorts to random exploration. Evaluation shows the significant improvement in planning efficiency.
For higher-dimensional systems, the memory requirements of the roadmap can be improved by considering sparse representations [36]. Furthermore, learned reachability estimators can assist in efficient roadmap construction and online queries. This work assumes an accurate model of the environment and the robot, which complicates deployment on real systems. This motivates integrating the proposed motion planner with system identification, state estimation and feedback control to track the planned trajectory.